value 2.00 points A random sample of 16 pharmacy customers showed the waiting times below in minutes) 1 20 12 17 2 20 2717 24 18 1517 16 17 13 20 20 匯clickhere for the Excel Data File Find a 90 percent confidence interval for, assuming that the sample is from a normal population. (Round your standard deviation answer to 4 decimal places and t-value to 3 decimal places. Round your answers to 3 decimal places.) to he 90% confidence interval
Answer #1

The 90% confidence ittnerval for population mean is given by

\overline{x}-E< \mu <\overline{x}+E

Where,

E = tc*\frac{s}{\sqrt{n}}

From data we get,

n = 16, \overline{x} = 18.38, s = 3.8101

c = 0.90, \alpha =1-c=1-0.90=0.10

df = n-1 = 16-1 = 15, tc=t_{\alpha /2,n-1}=t_{0.10/2,15}=1.753

So,

E = 1.753*\frac{3.8101}{\sqrt{16}}

= 1.669

Hence the confidence interval is given by

18.38-1.669 < \mu <18.38+1.669

16.711 < \mu <20.049

The 90% confidence interval 16.711 to 20.049.